OpenStudy (anonymous):

what is the trigonometric form of -4

OpenStudy (anonymous):

\[Let -4=r \left( \cos \theta+\iota \sin \theta \right) \] equating real and imaginary parts \[r \cos \theta=-4,r \sin \theta=0\] square and add \[r ^{2}=16+0=16,r=4\] \[4\cos \theta=-4,\cos \theta=-1=\cos \left( 2n+1 \right)\pi,where n is an integer.\theta=\left( 2n+1 \right)\pi \] for n=0 \[\theta=\pi \] \[-4=4\left( \cos \pi+\iota \sin \pi \right) \]